1 Distinguish noise from musical note.

2. The length of air column at which the first resonance was observed when a vibrating tuning fork was placed at the mouth of a resonance tube is 20 cm. Calculate

(i) The wavelength of the air column,

(ii) The frequency of the fork (assume the speed of sound in air is 340 m/s)

(iii) What is the frequency of the fork that would give resonance at 30 cm of the tube?

3. What is frequency and how is it related to the pitch of a note.

4. State two ways by which the pitch of a note by the string of a guitar can be changed.

5. What is meant by resonance

6. Explain a practical example where resonance is disadvantageous, and show how it can be avoided.

7. In a resonance tube experiment, a vibrating fork of frequency 256 Hz gave resonance when the water level was 30.0 cm below the open end of the tube. If the next position of resonance was 100 cm, what is the velocity of sound in air?
1
Expert's answer
2020-06-03T12:07:05-0400
  1. Noise is a mix of many different frequencies, musical note is just one frequency.

2. (i) The wavelength of the air column is


"\lambda=4l=4\cdot0.2=0.8\text{ m}."

(ii) The frequency of the fork (assume the speed of sound in air is 340 m/s)


"f=v/\lambda=340/0.8=425\text{ Hz}."

(iii) What is the frequency of the fork that would give resonance at 30 cm of the tube?


"f_1=f\frac{l}{l_1}=425\frac{20}{30}=283\text{ Hz}."

3. The frequency is how many times per second molecules of a medium oscillate, higher frequency corresponds to higher pitch. Low frequency is low pitch.

4. By changing tension or length of the string.

5. Rapid increase of amplitude of oscillations when the frequency of the external oscillating force approaches or equal to the natural frequency of the system.

6. It is disadvantageous in bridges. It can be avoided by changing the span or stiffness of the bridge, or by controlling the traffic flow through the bridge.

7. The velocity of sound in air is


"v=2f(L_2-L_1)=358\text{ m/s}."